1562 ____ ____ ____ ____ ____ 1874
I'm in 4th grade, can someone tell me the way to find these answers.

Answer by edjones(7311)

1874-1562=312
312/6=52
1562+52=1614+52=1666+52=1718+52=1770+52=1822+52=1874
.
Ed